# Vogelpothsweg 87, D-44227 Dortmund, Germany

### v1

@@ 1 Introduction
The design of a finite element scheme that ensures preservation of local and/or local bounds for numerical solutions to a hyperbolic problem typically involves construction and local adjustment of a dissipative shock capturing term. In most cases, a finite element counterpart of the local Lax–Friedrichs (Rusanov) method is corrected by adding limited antidiffusive terms. In edge-based schemes, the additional terms consist of fluxes between pairs of neighboring nodes guermond2018 ; convex ; fctbook ; kuzmin2010 ; luo1994 . In element-based alternatives, diffusive and antidiffusive correction terms are assembled from element vectors that possess the zero sum property dobrev2018 ; preprint ; kuzmin2020 ; lohmann2017 . A comprehensive review of the state of the art can be found in the recent book book .
@@ 5 Concluding remarks
Finally, we remark that the low-order scheme presented in Section 2 is equivalent to a subface finite volume (FV) discretization FV on a dual mesh $\mathcal{T}_{h}^{*}$ of control volumes associated with the nodal points $x_{i}$ of a finite element mesh $\mathcal{T}_{h}$. In the FV version, vertices become cell centers, while cell centers become vertices (cf. selmin1993 ; selmin1996 ). This observation makes it possible to incorporate convex limiting of FCT and MCL type into FV schemes for general polygonal and polyhedral meshes. The virtual finite element MCL method proposed by Abgrall et al. pampa is also applicable to such meshes, and its derivation is based on the RD design philosophy.

### v2

@@ 1 Introduction
The design of a finite element scheme that ensures preservation of local and/or global bounds for numerical solutions to a hyperbolic problem typically involves construction and local adjustment of a dissipative shock capturing term. In most cases, a finite element counterpart of the local Lax–Friedrichs (Rusanov) method is corrected by adding limited antidiffusive terms. In edge-based schemes, the additional terms consist of fluxes between pairs of neighboring nodes guermond2018 ; convex ; fctbook ; kuzmin2010 ; luo1994 . In element-based alternatives, diffusive and antidiffusive correction terms are assembled from element vectors that possess the zero sum property dobrev2018 ; preprint ; kuzmin2020 ; lohmann2017 . A comprehensive review of the state of the art can be found in the recent book book .
@@ 5 Concluding remarks
Finally, we remark that the low-order scheme presented in Section 2 is equivalent to a subface finite volume (FV) discretization FV on a dual mesh $\mathcal{T}_{h}^{*}$ of control volumes associated with the nodal points $\mathbf{x}_{i}$ of a finite element mesh $\mathcal{T}_{h}$. In the FV version, vertices become cell centers, while cell centers become vertices (cf. selmin1993 ; selmin1996 ). This observation makes it possible to incorporate convex limiting of FCT and MCL type into FV schemes for general polygonal and polyhedral meshes. The virtual finite element MCL method proposed by Abgrall et al. pampa is also applicable to such meshes, and its derivation is based on the RD design philosophy.